Oxford Materials IT Committee Meeting


A meeting of the Department of Materials IT Committee was held at 11am on Tuesday 10th May in Banbury Road Committee Room.


  • 1 Apologies for absence and Membership of the Committee
  • 2 Matters outstanding from the previous meeting
  • 3 Matters for discussion (attached)
  • 4 Head of IT report (attached)
  • 5 Finances
  • 6 AOB
  • 7 Date and time of next meeting:  Tuesday 25th October 2016 at 11am in 21 Banbury Road Committee Room 


See minutes below.


Update on recent infrastructure outages (attached document)

Verbal reports on the status of various projects:

IT services projects:  TONE, Chorus, OpenAccess
Departmental projects: Photocopiers, GreenImpact, iLab
Information from other Committees: MPLS ICT Panel, Divisional Software Licensing Group


The report below summarises activities since the last meeting of ITC, as categorised by topic. The report is intended for record rather than for discussion at ITC meeting although queries may arise.

Personnel: Re-advertising the Grade 6 IT staff position is still on hold pending the outcome of a wider Review of IT. After numerous delays, IT services have provisionally scheduled the Review of IT for June. The office space formerly occupied by that staff position has been cleared and reallocated as a research laboratory.

Admin Network New computer for Tom Lake who joined the department as HR manager. Tom Lake added to FilemakerPro using last spare license. Swapped new desktop for Barry Fellows. Swapped failed desktop on visitor door entry system. Swapped failed desktop for Karen Disney.

Teaching Laboratory Undergraduates continue to request print quota increases due to printing lecture notes. No other changes.

Digital Signage: Electronic display signage is now operational in HR reception and Begbroke Hirsch entrance. Lorraine Laird updates the content. The system could be expanded to have more display boards in other buildings if requested. 

Lecture theatres / meeting rooms There are still plans for Skype functionality in ETBCR and BRCR but this will be rolled into wider lecture theatre upgrades identified as a funding priority for summer (several users have complained about the low brightness of the existing old data projectors). PJW has implemented a proof of concept wireless presentation system for testing by lecturers before rollout around the department. QIP meeting room upgraded with wireless presentation specified by QIP members.

Printers and photocopiers: There have been problems with the stability of the print/photocopy system. Plans to decommission the main admin colour printer and teaching lab colour printer (moving print jobs to the photocopiers) have been postponed until the summer.

Digital Print-room The light bench for photography of specimens has been disposed due to lack of use since the photographic unit closed in 2008.

Library The self-issue system has been upgraded.

MML Four replacement motherboards were purchased after the January aircon failure. Ironman continues to have issues. Aircon partially failed again on 10th April and is still running at risk.

FIM Still need to replace image acquisition PC on old 3DAP. 

EM The delayed deletion of user data from EM storage array is now becoming a priority due to lack of disk space.

MFFP Monthly patching and rebooting of 13 systems and storage arrays. MFFP website is desperate for upgrading due to Drupal6 end of life in April. New computer system (512Gb RAM) was purchased to support new Xray tomography unit. New storage array purchased to backup existing array (disaster failover tested but need to test again).

OMCS Three new UPS systems purchased and due to be commissioned. New user added to ACT.

PGB Salto door entry system still needs fixing to work with university cards rather than separate cards. Steady stream of new people and equipment.  Borri UPS underwent annual maintenance.

Website Daily/weekly editing of department website and news items. New 64bit webserver is still being developed to replace end of life old 32bit webserver. CmsMadeSimple need upgrading 1.13 to 2.1 on new server. MFFP website needs migrating to Drupal7 since Drupal6 is almost end of life. All websites need migrating to new webserver and configuring for https.

Publications Weekly editing of Departmental Newsletter. 13 sets of business cards. Printed final version of Graduate handbook.  Signage for March opendays.


As part of the university TONE and Odin project, feeding a new fibre into Hume-Rothery building was not possible due to the discovery of asbestos on the main (and only) route to the data cabinet. Unfortunately, this means that no new network connections can be made within the Hume-Rothery building (or until a new route is created). As a workaround some unused fibres in the existing cable have been terminated and on 8th May the department’s backbone connection will be migrated to the new FroDo units connected to the new university backbone.
As part of the university TONE and Odin project, a new FroDo cabinet has been installed and new network feeds run to the existing data cabinets (£3k). Most of the network switches in BR have been replaced with new gigabit Extreme switches.

Security Oxcert have released 30 security bulletins. On reported Dridex infection resolved by reinstallation. Several minor virus infections dis-infected. One remote access infection security block).

WDE Still only six user’ devices encrypted – need to expand uptake targeting senior staff. Limited by time/manpower….

Disposal A large load of WEEE was transferred from Begbroke to central site but the actual disposal van was cancelled in March due to time constraints.  A WEEE skip is currently being used to clear Richard Cripps’ office and dispose of other miscellaneous IT equipment.

Software No significant changes. We are able to deploy Windows 10 but still prefer to deploy Windows 7. Microsoft are pushing hard for home and business users to migrate – deadline for free home upgrade end of July.

Disaster Recovery and Backup Several disk fails have occurred and data has been recovered or restored from backups.  Partial hardware fail on main Symmetra UPS has reduced redundancy but is manageable in the medium term (second-hand replacement power module being sought <£1200). Purchased 12 UPS systems from closing-down UPS supplier at a good price (£300 for 3kVA), mostly for research groups. This has turned into a big time-consuming project to specify, purchase, setup network monitoring and configure auto-shutdown of all UPS protected devices – still need to roll all these UPS’s into production.

Documentation updates Documentation on the private Mediawiki continues to expand, but time for documentation has been limited due to being short staffed.

New people & accounts February 3; March 10; April 4;

New Device registrations February 10 ; March 18; April 10;


A meeting of the Department of Materials IT Committee was held at 11am on Tuesday 2nd February 2016 in the PR Meeting Room.
The following people attended: APD, PJW, RS, CA, RT, SPF, Flynn Castles, Andy Naylor, Lewys Jones, Jiale Wang

1. Apologies for Absence and Membership of the Committee

Paul Bagot and Lewys Jones sent their apologies.

Professor Sergio Lozano-Perez (SLP) was welcomed to the committee as a faculty representative.

Alana Davies noted that this is to be her last meeting of ITC. The newly appointed Department Administrator, Charlotte Sweeney, is due to start in August. PJW thanked APD for her contributions to the committee and to the whole department since January 2006 and wished her well for the future.

 2. Matters arising from the minutes

(Items marked with * were not for discussion at the meeting)

a) *Draft Information Security Policy (Minute 2b)

Again there has been little opportunity to progress towards formalising the departmental Information Security Policy for approval by the ITC and subsequently DC. However PJW noted that the ISP documentation and toolkit has evolved considerably over the last few months and PJW is keen to revisit implementing a departmental IS policy, supported by PSG. There has been steady progress towards getting staff to complete the awareness module.

Action: PJW/PSG

b) *Office2013 for Admin (Minute 2c)

APD asked whether Office2013 or Office2016 could be rolled out more widely. PJW to investigate compatibility with central university systems such as macros in R12 reports.

Action: PJW

3 Matters for Discussion

a) Report on recent infrastructure outages.

PJW circulated a three-page report on the IT infrastructure instabilities and outages in Trinity Term 2016. The report presented a detailed timeline and technical description of the series of outages affecting the main data storage system behind the department’s virtual infrastructure.

Services affected by these outages included:

  • The main fileserver supporting all administrative and teaching functions, the commercial activities of OMCS, plus some research groups.
  • All departmental printing and photocopying systems 
  • Locally hosted websites (all except the department website) including room booking, car parking, equipment booking and numerous research group websites. 
  • Wireless provision just in the Rex Richards and Hume-Rothery buildings.
  • Email archives, containing all recent Administrator and Graduate Admissions emails.

APD noted that the whole of the department’s administrative systems were without access to files for 3.5 days (3/4/5/6 May) causing major disruption. These disruptions as well as being inconvenient will have caused financial loss, both in terms of time and also in terms of missing deadlines, such as grant applications.

After questions and discussion APD requested that IT staff put together a document detailing the main areas of risk associated with single points of failure within departmental IT systems, and presenting costed proposals on how to minimize future risks. This document could then feed into the forthcoming Review of IT by IT services.


b) Verbal update on various projects

PJW updated the committee verbally upon the status of the following projects:

  • TONE – replacement backbone network. http://projects.it.ox.ac.uk/tone/
    Department of Materials was successfully migrated to the new university backbone on Monday 9th May. The changeover was achieved with minimal disruption (65 seconds of network outage for the whole department, a few minutes for Rex Richards building and a couple of minute outages for 14 Parks Road).

  • Chorus – the new telephone system http://projects.it.ox.ac.uk/icp/
    Also see official Chorus newsletters http://projects.it.ox.ac.uk/icp/newsletter
    Another pilot project for 250 users is currently running. The new system is scheduled to be rolled-out to all users by the end of 2017. Units are to be given at least 3 months’ notice prior to migration.

  • OpenAccess http://openaccess.ox.ac.uk/
    The new HEFCE rules came into force on 1st April. For every article published authors need to Act on Acceptance to ensure the article is eligible for REF 2021. Existing subject repositories (such as ArXiv) are not compliant for REF. Patrick Grant is Open Access coordinator and Paul Warren is Open Access monitor, and together they will liaise with researchers to ensure that all departmental publications are submitted as required.

  • Java Runtime Environment Upgrade
    R12 financials migrated to JRE 1.8 in February so all core systems are now compatible with supported secure versions of Java.

  • Information Security Improvement Programme (ISIP)
    This department underwent an external webserver security audit as part of the ISIP project. PJW will move forward with some recommendations while commissioning a new department webserver. This department still needs to review and adopt an Information Security Policy.(Minute 2a). There is now a new Information Security website providing clearer guidance to users. https://infosec.ox.ac.uk. There has been a steady increase in the number of staff members who have completed the information awareness module and PJW continues to monitor progress and target people.

Departmental projects:

  • Photocopiers - The copiers are being used for printing and scanning. APD’s copier continues to behave erratically and needs attention. 

PJW reported a recent issue with label printing on photocopiers requiring manual feed of labels which resulted in numerous users releasing their print jobs but not getting the paper output until the manual feed request had been cleared, at which point all the released jobs were printed together – with the potential for sensitive data being viewed. It was agreed to highlight this in a future Mildred article to the newsletter.

Action: PJW

In Hilary Term there was also a recommendation to implement monthly email reports of print/copy usage to all users showing the costs in order to reduce usage.

Action: PJW

IT staff are currently resolving technical details over implementing restricting printing from the unrestricted Teaching Laboratory microscope computers, prior to removing the colour printer and making the undergraduates print to the photocopier which uses a print-on-release method so there should not be any uncollected print jobs.

Action: PJW/RS

  • iLab - APD has been talking with iLab about further refinement of reporting and the need for collecting actual usage data and implementing quotas. The new XRD equipment will be incorporated into the EM core, rather than having a separate core.
  • Green Impact – the team submitted the annual workbook in April and the department has been audited. We currently await the outcome.

Information from other committees:

  • MPLS ICT panel – the Trinity Term meeting yesterday received a status report from IT Services (circulated via email). The main topics were a presentation about the difficulties of backing-up large storage, and a summary of MPLS HoD’s views on research IT capital investment priorities (sync&share dropbox, ARC expansion, systems to support statutory compliance, systems to support research grant management).
  • Divisional Software Licensing Group http://help.it.ox.ac.uk/sls/background
    Origin site licence is due for renewal in July (managed by PJW). Microsoft license costs are expected to increase by 20% before June. Matlab site license renewal in August is intended to be funded from divisions on the basis of pro-rate usage data.

4. Head of IT Report
A written report was circulated with the agenda.

5. Finances
The areas strategically identified as needing attention (see minutes of the last ITC meeting) were incorporated into proposals for the 2016/2017 budget planning meeting, which was approved. Details circulated via email. Notably TMcA agreed that the university’s Origin site license renewal managed by PJW can be processed through the department accounts. PJW needs to review current expenditure and submit all stock journal requests prior to the financial year end on 31st July.

Action: PJW

 6. AOB

SLP requested that the departmental display signage system be extended into the Holder Building cafeteria in order to counter-balance the Engineering display signage.

Action: PJW

 RT noted that the visitor information leaflet needs some Begbroke information updating, including the building site map.

Action: PJW

7. Date of Next Meeting

The next meeting is Tuesday 25th October 2016 at 11am in 21 Banbury Road Committee Room .

The meeting closed at 12.40pm.